hallm:

Solve for x. 56x=15 Enter your answer in the box. x =

hallm:

please help i dont know how to get the answer

Karla:

Divide the equation by 56 from both side in order to separate x it will be 56/56 x=15/56 56/56 =1. While 15/56=0.27 So 1x=15/56 Or you can say x=15/56 or 0.27